OpenStudy (anonymous):

Just wanted to see if this was right. Solve in quadratic equation. x^2 + 3x = 10 x^2 + 3x - 10 = 10 - 10 x^2 + 3x - 10 = 0 x + 5x - 2 = 0 x + 5 = 0 or x - 2 = 0 x = -5 or x = 2 If it's not please say where the mistake was made. Thanks so much.

OpenStudy (anonymous):

x^2 +3x -10 = 0 x^2 +5x -2x -10 = 0 x(x+5) -2(x+5) = 0

OpenStudy (anonymous):

that is correct.

OpenStudy (anonymous):

x=-5 and x=2

OpenStudy (anonymous):

Ok, Don't get it. That just gave me equation it gave me. But I needed two different answers for what x= ?. I am. yay. Was all kinds of confused for a moment there. Thank you.

OpenStudy (anonymous):

ok, x (x+3) = 10 (x-2) (x+5) = 0 so x=2 and x=-5

OpenStudy (jhannybean):

But why would you write `x(x+3)=10` @felavin ?

OpenStudy (jhannybean):

Your first 3 steps seem to be fine, @bitesize007 , then i'm not sure how you got to x+5x-2=0

OpenStudy (anonymous):

we have x^2 + 3x = 10 so x (x+3) = 10 and then: (x-2) (x+5) = 0 so x=2 and x=-5

OpenStudy (anonymous):

this is the fast way @Jhannybean

OpenStudy (anonymous):

Should the 3 have stayed? I don't really know how it got turned into a 5 to be honest. Saw it down in the book to almost the same equation so just copied what it did in the book bu steps.

OpenStudy (jhannybean):

Just according to what you have written, up to the third step: `x^2 +3x-10=0` The 4th step you have written there looks wrong to me: `x+5x-2=0` From the third step that you have there, I would go straight to factoring, that is, find two numbers that multiply to give `-10` and add to give `+3`

OpenStudy (jhannybean):

SO you have everything right, just remove that 4th step: `x+5x-2=0`you have written there.

OpenStudy (jhannybean):

Are you good? :)

OpenStudy (anonymous):

Just remove the 4th step or remove all from there and rewrite it like this ? x^2 + 3x - 2 = 0 x + 3 = 0 or x - 2= 0 x = -3 0r x = 2

OpenStudy (jhannybean):

Any case, just remove the 4th step you have written there and you will be fine.

OpenStudy (anonymous):

Alright, Thanks for your help.